1. Estimate the volume of the earth in m3 by assuming that its radius is 6400km.



2. A car accelerates from zero to 20 m/s in 5 s.



i. Calculate the acceleration of the car.


ii. If acceleration is assumed to be constant, how far


will the car travel in 15 s?


iii. Calculate the speed of the car after 15 s.



3. A KTM Commuter train travel between two station, A and B. The train from rest from station A and accelerates at a constant rate for 0.5 k.m until it reaches a speed of 30 m/s. It then travel at this constant speed for Ts before decelerating at a constant rate of 1.5 m/s and coming to rest at station B. Determine the value of T if the distance between the two stations 6km. Calculate also the total time taken for the whole journey.

1
Expert's answer
2022-01-24T08:45:11-0500

1.

"V=4/3\pi R^3\\
=4/3*3.14*(6400\:\rm km)^3=1.1*10^{12}\: km^3"


2.

(i)

"a=\frac{\Delta v}{\Delta t}=\frac{20}{5}=4\:\rm m/s^2"

(ii)

"d=at^2/2=4*15^2/2=450\:\rm m"

(iii)

"v=at=4*15=60\:\rm m/s"